OpenStudy (anonymous):

An airplane took 3 hours to fly 1200 miles against a headwind. The return trip with the wind took 2 hours. Find the speed of the plane in still air and the speed of the wind.

OpenStudy (anonymous):

let rate of plane in still air be x let rate of wind be y (x-y)*4 =1200 (x+y)*3=1200 4x-4y=1200 3x+3y=1200 3(4x-4y)+4(3x+3y)=3*1200+4*1200 12x-12y + 12x +12y = 8400 24x=8400 x=350 mph rate of plane in still air (x+y)*3=1200 (350+y)*3 =1200 1050+3y=1200 3y=1200-1050 3y= 150 y=50 mph rate of wind